How many 5 digit munber can be made by using 1,2,3,4,5,6,7,8 in the number are greater 4321 if repetation js not allowed

Grade:11

1 Answers

Arun
6 years ago
Case 1-
 
If 1st digit from left is greater than 4
Hence
Total numbers = 4 *7*6*5*4 = 1680
 
Case 2-
 
First digit from left is 4
Then option for second digit , this also makes 2 option as 
It is 3 then also for third digit there are 2 options as it will be 2 or greater than 2.
Hence total numbers in 2nd case
= 1* 1*1*4 + 1* 1* 4* 5 + 1* 4* 6*5 
= 4 + 20 + 120 = 144
 
Hence total numbers = 1680 + 144 = 1828
 
Regards
Arun (askIITians forum expert)